OpenStudy (anonymous):

If 10,800 cm^2 of material is available to make a box with a square base and an open top, find the largest possible volume of the box.

OpenStudy (anonymous):

well since the top is opend we have a cube with 5 sides. Since all the sides of a cube are the same the area of one side is 5x=10,8 x=2.16 now we can calculate the side of the box(Cube) and it is a^2=2.16 a=1.46 so now the volume of the cube is a^3=(1,46)^3=3.11 cm^2

OpenStudy (anonymous):

That doesn't make sense, as the total available amount of material is in the thousands. Your answer, however, is approximately 3.0.

OpenStudy (anonymous):

I tought that the area is 10(point)8, sorry for the mistake, just follow my logic and replace it with the big number

OpenStudy (anonymous):

So the answer is 3110?

OpenStudy (anonymous):

5x=10800 x=2160 a^2=2160 a=46.47 a^3=100350 cm^3

OpenStudy (anonymous):

I'm afraid that isnt correct. This is a calculus question

OpenStudy (anonymous):

the volume is 108 000cm^3